4. What speed would a 2,000-kg sport utility vehicle have to attain in order to have i) the same momentum? ii) the same kinetic energy?


1
Expert's answer
2021-09-28T11:13:29-0400

Assume the momentum is p, the energy is E. Then


"v=\frac pm=\frac p{2000},\\\space\\
v=\sqrt{\frac{2E}{m}}=\sqrt{\frac{E}{1000}}."

Plug in the values of p and E and get the speed.